Find 3 consecutive integers such that 5 less than 3 times the smallest is the same as 21 less than the largest.
Answer by oberobic(2304)

Three consecutive integers:
x
x+1
x+2
.
3x -5 = x+2 -21
2x -5 = -19
2x = -14
x = -7
.
x+1 = -6
x+2 = -5
.
Does that solve the problem?
3*(-7) -5 = -26
-5-21 = -26
Correct.
.
Answer: The three consecutive integers are: -7, -6, and -5.
.
Done.